Bill Phipps is 6 times as old as his 5-year old niece. How long will it be before he is twice as old as she?
Thanks for your help :)

Answer by jim_thompson5910(35256)

Bill Phipps is 6 times as old as his 5-year old niece

So Bill is 6*5 = 30 years old

In x years, he will be 30+x years old

In x years, she will be 5+x years old

If you want to know when he will be twice her age, simply say

30+x = 2(5+x)

Then solve for x

30+x=2(5+x)

30+x=10+2x

x=10+2x-30

x-2x=10-30

-x=10-30

-x=-20

x=(-20)/(-1)

x=20

So it will take 20 years for him to be twice as old as her.

Notice how adding 20 years to 30 makes him now 50. Adding the same 20 years to 5 makes her now 25. Finally, notice that 50 is double of 25. So the answer checks.